OpenStudy (anonymous):

6x-3y=21 2x+2y=22 elimination by multiplacation.

OpenStudy (mertsj):

Multiply the bottom equation by -3 and add the result to the top equation.

OpenStudy (anonymous):

-6X+6y=-66

OpenStudy (anonymous):

Well what do you mean by eliminate by multiplication? What I did was I set it up so that you have: 6x-3y=21 -3(2x+2y=22) this will give you 6x-3y=21 and -6x-6y=-66 This system enables you to solve for y. -3y-6y=21-61 -9y=-45 y=5 Now, you can just plug y back in to any of the 2 equations given. x=6

OpenStudy (anonymous):

thank you!